### v1

# Scaling Laws Meet Model Architecture: Toward Inference-Efficient LLMs

Abstract
Scaling the number of parameters and the size of training data has proven to be an effective strategy for improving large language model (LLM) performance. Yet, as these models grow increasingly powerful and widely deployed, the cost of inference has become a pressing concern. Despite its importance, the trade-off between model accuracy and inference efficiency remains underexplored. In this work, we examine how key architectural factors, hidden size, the allocation of parameters between MLP and attention (mlp-to-attention ratio), and grouped-query attention (GQA), influence both inference cost and accuracy. We introduce a conditional scaling law that augments the Chinchilla framework with architectural information, along with a search framework for identifying architectures that are simultaneously inference-efficient and accurate. To validate our approach, we train more than 200 models spanning 80M to 3B parameters and 8B to 100B training tokens, and fit the proposed conditional scaling law. Our results show that the conditional scaling law reliably predicts optimal architectural choices and that the resulting models outperform existing open-source baselines. Under the same training budget, optimized architectures achieve up to 2.1% higher accuracy and 42% greater inference throughput compared to LLaMA-3.2.
1 Introduction
Scaling law studies Kaplan et al. (2020); Hoffmann et al. (2022); Muennighoff et al. (2023); Krajewski et al. (2024); Abnar et al. (2025) have shown that increasing model parameters, training tokens, dataset quality, and compute budget consistently reduces pre-training loss, improves downstream task performance Hendrycks et al. (2021); Austin et al. (2021), and enables the emergence of novel capabilities Wei et al. (2022). These insights have driven the development of many state-of-the-art large language models Touvron et al. (2023); Yang et al. (2025); Guo et al. (2025).
However, as the field advances, it has become increasingly clear that focusing exclusively on training overlooks the practical challenges of deploying these models at scale Chien et al. (2023); Wu et al. (2024); Muhamed et al. (2023). A major limitation of existing scaling laws is their omission of inference costs, which constitute the dominant expense in deploying large models in real-world applications Sardana et al. (2023); Park et al. (2024). Moreover, the growing use of LLMs in reasoning systems highlights the need for scaling laws that account for inference costs Snell et al. (2024); Brown et al. (2024); Luo et al. (2024); Qi et al. (2024); Guan et al. (2025). Therefore, we ask the following question:
Can we explicitly capture the trade-off between inference efficiency and accuracy of large language models?
To address this question, a recent study Sardana et al. (2023) proposed scaling laws that incorporate the total FLOPs from both training and inference. However, their formulation requires estimating the total number of tokens generated over a model’s entire lifespan. Because inference is performed repeatedly during deployment, this assumption renders the proposed scaling law impractical for real-world use. Another study Bian et al. (2025) extends Chinchilla scaling laws by incorporating model architecture. However, this work has notable limitations. First, the study considers only the aspect ratio, defined as hidden size over number of layers, as the architectural factor. Yet, as shown in Figure 1, aspect ratio alone fails to capture the full range of factors that influence inference efficiency in large language models. Second, the depth of the model strongly influences accuracy: cutting layers tends to impair the model’s generalization after fine-tuning Petty et al. (2023). Finally, the study lacks a general framework for incorporating broader architectural factors, including hidden size and GQA, into scaling laws.
Figure 1: Although larger models generally achieve lower inference throughput than smaller ones, Qwen2.5-1.5B outperforms Qwen3-0.6B. Despite having the same number of layers, Qwen2.5-1.5B benefits from a higher hidden size, GQA, and mlp-to-attention ratio.
In this work, we fix the number of layers and study the effect of other architectural factors, including GQA, hidden size, and the mlp-to-attention ratio. This design choice is motivated by recent open-weight models such as LLaMA Touvron et al. (2023), Qwen Yang et al. (2025), Gemma Team et al. (2024a), and Phi Abdin et al. (2024), which, despite having a comparable number of parameters, adopt markedly different architectural designs.
Our primary goal is to investigate how model architecture influences both inference efficiency and model accuracy. We begin by comparing the inference efficiency of models with identical parameter counts but varying architectures. Next, we train over 200 models, ranging from 80M to 297M parameters on up to 30B tokens, to systematically characterize the relationship between architectural design and accuracy. Guided by these empirical findings, we introduce a conditional extension of the Chinchilla scaling laws that incorporates architectural parameters, establishing a general framework for identifying model architectures that balance inference efficiency and performance.
Finally, we validate this framework by fitting the proposed scaling law on models between 80M and 297M parameters, and evaluating its predictions when scaling up to 3B-parameter models. Our results demonstrate that, under identical training setups, the derived optimal 3B-parameter architecture achieves $42\%$ higher inference throughput than the LLaMA-3.2-3B architecture, while maintaining better accuracy.
2 Background
Accurately predicting the performance of large language models during scaling is essential. This enables us to answer key questions: (i) what is the optimal allocation of available resources between model size and training tokens, and (ii) what performance gains can be expected from additional resources? Fortunately, the model loss has been observed to follow a power-law relationship with respect to the number of parameters $N$ and training tokens $D$ Hoffmann et al. (2022); Muennighoff et al. (2023) with:
$L(N,D)=E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}$ (1)
where $L$ is the model loss, $N$ is the number of total parameters and $D$ is the number of tokens used for training and $A$, $B$, $E$, $\alpha$, $\beta$ are parameters to be learned.
To fit the learnable parameters in Eq. (1), Chinchilla Hoffmann et al. (2022) employs two strategies: (i) training models with a fixed number of parameters while varying the number of training tokens, and (ii) training models under a fixed compute budget [^1] The compute cost is approximated as $\text{FLOPs}(N,D)\approx 6ND$ in Hoffmann et al. (2022); Muennighoff et al. (2023), where $N$ denotes the number of parameters and $D$ the number of training tokens. In this work, we adopt the same settings as prior studies., varying both parameters and tokens. The resulting data are combined to fit the learned parameters in Eq. (1). With the fitted scaling laws, Chinchilla addresses the following question to determine optimal allocation:
$\arg\min_{N,D}L(N,D)\text{ s.t. }\text{FLOPs}(N,D)=C$ (2)
where $C$ denotes the resource constraint, $N$ the total number of parameters, and $D$ the number of training tokens.
In this paper, we do not address how to optimally allocate compute between model size and training data under a fixed compute budget. Instead, our focus is on identifying model architectures that optimize inference efficiency and accuracy under fixed parameter and token budgets. For example, given a model with 7B parameters trained on 14T tokens, we study how to design an architecture that satisfies both efficiency and accuracy requirements.
3 Model Architecture-Aware Scaling Laws
3.1 Model Architecture Variations
The architecture of a decoder-only transformer is composed of a sequence of stacked decoder blocks, each sharing the same structure to facilitate model-parallel deployment across devices. Under this design, the overall architecture of dense LLMs is primarily determined by the hidden size and the MLP intermediate size, which together specify the attention and MLP layers structure. This work studies the optimal model architecture given a fixed total number of non-embedding parameters $N_{\text{non-embed}}$ (at different levels). Although the number of layers $n_{\text{layer}}$ also plays a critical role (closely related to aspect ratio (Petty et al., 2023)), varying $n_{\text{layer}}$ under a fixed $N_{\text{non-embed}}$ substantially impacts both inference cost and accuracy (Tay et al., 2021; Alabdulmohsin et al., 2023). Therefore, we fix $n_{\text{layer}}$ and focus on the effects of hidden size $d_{\text{model}}$ and the mlp-to-attention ratio $r_{\text{mlp}/\text{attn}}$ on inference efficiency (§3.2) and accuracy (§3.3), noting that $n_{\text{layer}}$ still varies across different $N_{\text{non-embed}}$ levels. In §3.3, we introduce a conditional scaling law to predict the performance of architectural variants, and in §3.4, we present a lightweight framework for identifying architectures that optimally balance inference efficiency and accuracy.
Note that the number of attention parameters is primarily determined by the hidden size $d_{\text{model}}$ and the attention projection dimension, since most open-weight models adopt non-square $q,k,v$ projection matrices, as seen in Gemma (Team et al., 2024a) and Qwen3 (Yang et al., 2025). For consistency, we fix the per-head dimension $d_{\text{head}}$ to 64 for models with $N_{\text{non-embed}}$ $\leq$1B and to 128 for models with $N_{\text{non-embed}}$ $\geq$3B. Consequently, to maintain a constant $r_{\text{mlp}/\text{attn}}$, we adjust the number of attention heads $n_{\text{head}}$ rather than altering the projection dimension directly. This design choice also provides flexibility to incorporate architectural variants such as grouped-query attention.
Figure 2: Inference throughput vs (left) hidden size $d=d_{\text{model}}$ and (right) mlp-to-attention ratio $r=r_{\text{mlp}/\text{attn}}$ on the 8B model. Under a fixed parameter budget $N_{\text{non-embed}}$, larger hidden sizes and higher mlp-to-attention ratios improve inference throughput for varying batch sizes.
3.2 Inference Efficiency
Inspired by the success and widespread adoption of open-weight dense models such as Qwen3 (Yang et al., 2025), LLaMA-3.2 (Dubey et al., 2024), and the Gemma-2 (Team et al., 2024b) family, we construct architectural variants by modifying the configurations of the LLaMA-3.2 and Qwen3 dense models (Figure 11-13 in Appendix E). In addition to hidden size and the mlp-to-attention ratio, we find that group-query attention has a critical impact on inference efficiency, even though it only modestly reduces the number of attention parameters (by shrinking the key and value matrices). To disentangle these effects, we conduct controlled ablations of hidden size, MLP-to-attention ratio, and GQA under the following setups:
  • hidden size $d_{\text{model}}$: fix $N_{\text{non-embed}}$, $r_{\text{mlp}/\text{attn}}$ and GQA=4=4, vary $d_{\text{model}}$ and number of attention heads $n_{\text{head}}$ (Figure 2 left).
  • mlp-to-attention ratio $r_{\text{mlp}/\text{attn}}$: fix $N_{\text{non-embed}}$, $d_{\text{model}}$ and GQA=4=4, vary $n_{\text{head}}$ and intermediate size (Figure 2 right).
  • GQA: fix $N_{\text{non-embed}}$, $d_{\text{model}}$ and $r_{\text{mlp}/\text{attn}}$, vary $n_{\text{head}}$ and number of key-value heads (Appendix E).
Figure 2 shows the ablation of varying hidden sizes $d_{\text{model}}$ and mlp-to-attention $r_{\text{mlp}/\text{attn}}$ on the LLaMA-3.1-8B model variants. We observe that larger hidden size (or fewer attention heads) and higher mlp-to-attention ratios improve inference throughput. Similar trends are observed in the LLaMA-3.2-1B and 3B model variants (Appendix E). These gains arise in part because larger $d_{\text{model}}$ and higher $r_{\text{mlp}/\text{attn}}$ reduce the total FLOPs, as detailed in the inference FLOPs analysis (Appendix H). In addition, these architectural choices shrink the KV cache, lowering I/O cost during inference and further improving throughput Adnan et al. (2024). Figure 10 in Appendix E presents the GQA ablation, confirming prior observations Ainslie et al. (2023) that increasing GQA consistently improves inference throughput. A comparable set of ablation experiments on Qwen3 models, also reported in Appendix E, further corroborates these findings.
3.3 A Conditional Scaling Law
Improving inference efficiency should not come at the expense of significantly reducing model accuracy, making it crucial to understand how architectural choices affect accuracy and training loss. Because training large-scale language models is prohibitively expensive, a common strategy is to study smaller models and use scaling laws to extrapolate insights to larger scales, for example, the Chinchilla scaling laws (Hoffmann et al., 2022). However, incorporating multiple architectural factors into such laws remains challenging. To address this, we examine the effect of architectural choices on training loss $L$ in a conditional manner, varying one factor at a time while keeping the others fixed.
hidden size $d_{\text{model}}$.
We note that $d_{\text{model}}$ generally scales linearly with $\sqrt{N_{\text{non-embed}}}$. Assuming squared attention weight matrices, the number of attention parameters $N_{\text{attn}}$ can be expressed as
$4d_{model}^{2}\propto N_{\text{attn}}=N_{\text{non-embed}}\times\frac{r}{r+1},$
where $r=r_{\text{mlp/attn}}$ is fixed, and the constant factor $4$ arises from the query, key, value, and output projection layers in each attention block. To capture this scaling behavior, we normalize $d_{\text{model}}$ by $\sqrt{N_{\text{non-embed}}}$ and examine its relation to loss $L$ in Figure 3. The resulting U-shaped curves $L(d/\sqrt{N}\mid r,N,D)$ exhibit nearly identical optima across different model sizes. Moreover, Figure 3 confirms that excessively large hidden sizes, which reduce the number of attention heads $n_{\text{head}}$, can degrade accuracy—a phenomenon consistently observed in prior analyses of transformer capacity and head allocation (Kaplan et al., 2020; Hoffmann et al., 2022).
Figure 3: Loss vs. hidden size: (Left) 80M model variants; (Center) 145M model variants; (Right) 297M model variants. Across model sizes, the relationship between training loss and $d_{\text{model}}/\sqrt{N}$ exhibits a consistent U-shaped curve when architectural factors such as GQA and the MLP-to-attention ratio are held fixed. The legend denotes the MLP-to-attention ratio $r=r_{\text{mlp}/\text{attn}}$ for each model.
Figure 4: Loss vs. MLP-to-attention ratio: (Left) 80M model variants; (Center) 145M model variants; (Right) 297M model variants. Across model sizes, the relationship between training loss and $r_{\text{mlp}/\text{attn}}$ exhibits a consistent U-shaped curve when architectural factors such as GQA and hidden size are held fixed. The legend denotes the hidden size $d=d_{\text{model}}$ for each model.
mlp-to-attention ratio $r_{\text{mlp}/\text{attn}}$.
Figure 4 illustrates how the loss varies with $r_{\text{mlp}/\text{attn}}$, conditioned on $d_{\text{model}}$ fixed at different levels, where we consistently observe a U-shaped curve $L(r\mid d/\sqrt{N},N,D)$. While the attention mechanism is central to the success of transformers (Vaswani, 2017), recent open-weight models have allocated a progressively smaller fraction of parameters to attention as overall model size increases (e.g., LLaMA and Qwen families). Our analysis indicates that this trend is not universally optimal: there exists an interior optimum in the allocation of attention parameters, and deviating from it in either direction degrades model performance. This suggests that careful tuning of the mlp-to-attention ratio is critical for scaling transformers effectively.
As shown in Figures 3 and 4, both hidden size and the MLP-to-attention ratio exhibit U-shaped relationships with training loss. To capture these trends, we fit the function $c_{0}+c_{1}\log x+c_{2}/x$ separately for $x=r_{\text{mlp/attn}}$ and $d_{\text{model}}/\sqrt{N_{\text{non-embed}}}$. This formulation effectively models the U-shaped behavior while ensuring sublinear growth as $x$ increases. However, incorporating $r_{\text{mlp}/\text{attn}}$, $d_{\text{model}}$, $N$, and $D$ into a unified, architecture-aware scaling law remains challenging. Since fitting a single all-purpose scaling law $L(d/\sqrt{N},r,N,D)$ is unrealistic across all possible configurations, we instead propose a two-step conditional approach:
  1. nosep
    For given $N$ and $D$, obtain the optimal loss $L_{\text{opt}}(N,D)=\min L(N,D)=\min\big(E+\tfrac{A}{N^{\alpha}}+\tfrac{B}{D^{\beta}}\big)$ from the Chinchilla scaling law (Eq. 1) as a reference point.
  2. nosep
    Calibrate the loss of architectural variants $L(d/\sqrt{N},r\mid N,D)$ relative to this reference.
We focus on two simple calibration schemes:
  • (multiplicative)
    $L(d/\sqrt{N},r\mid N,D)=(a_{0}+a_{1}\log(\frac{d}{\sqrt{N}})+a_{2}\frac{\sqrt{N}}{d})\cdot(b_{0}+b_{1}\log r+\frac{b_{2}}{r})\cdot L_{\text{opt}}$ (3)
  • (additive) $L(d/\sqrt{N},r\mid N,D)=(a_{0}+a_{1}\log(\frac{d}{\sqrt{N}})+a_{2}\frac{\sqrt{N}}{d})+(b_{1}\log r+\frac{b_{2}}{r})+L_{\text{opt}}$
Here, $a_{i}$ and $b_{i}$ are learnable parameters that are shared across all $N,D$. Unlike the unified formulation, the conditional scaling law assumes that the effects of $r_{\text{mlp}/\text{attn}}$ and $d_{\text{model}}$ on loss are separable. We further ablate joint, non-separable formulations in Appendix G, where we find that they yield inferior predictive performance.
3.4 Searching for Inference-Efficient Accurate Models
With the conditional scaling law, we can identify architectures that are both inference-efficient and accurate by solving the following optimization problem: given $N$, $D$, and a set of architectural choices $P$,
$\text{argmax}_{P}I_{N}(P),\hskip 20.00003pt\text{s.t.}\hskip 10.00002ptL(P\mid N,D)\leq L_{t},$ (4)
where $I_{N}(P)$ denotes the inference efficiency of an architecture $P$ with total $N_{\text{non-embed}}$ parameters, and $L_{t},(\geq L_{\text{opt}})$ is the maximum allowable training loss.
As shown in Figure 10 (Appendix E), GQA has a substantial impact on inference efficiency; However, unlike hidden size and the mlp-to-attention ratio, GQA does not exhibit a consistent relationship with loss (Figure 14) and is highly variable, making it challenging to identify settings that achieve both accuracy and efficiency. Fortunately, the search space for GQA is relatively small once $N_{\text{non-embed}}$, $d_{\text{model}}$, and $r_{\text{mlp}/\text{attn}}$ are fixed, since GQA must be a prime factor of the number of attention heads $n_{\text{head}}$. In practice, we perform a local GQA search by enumerating feasible values and applying early stopping once performance falls below that of the GQA=4=4 baseline. Algorithm 1 summarizes our overall framework for identifying inference-efficient and accurate architectures.
Input: Model parameters NN, training tokens $D$, target loss $L_{t}$; inference efficiency $I_{N}(\cdot)$; optional: the optimal loss $L_{\text{opt}}(N,D)$
Train smaller models to fit the Chinchilla scaling laws (Eq. 1) if $L_{\text{opt}}(N,D)$ is unavailable
Solve the constrained optimization (Eq. 4) for $d_{\text{model}}$$r_{\text{mlp}/\text{attn}}$ and corresponding architecture $P$
Perform a local search over GQA values with early stopping to maximize inference efficiency
return Final model architecture $\{P,\text{GQA}\}$
Algorithm 1 Searching for Inference-Efficient Accurate Model
4 Experiment Setup
We first detail the experimental setup of training, inference, and downstream task evaluation, and then describe how we derive the conditional scaling law and scale up to larger sizes.
Training Setup.
[…]
Inference Setup.
We evaluate the inference efficiency using the vLLM framework Kwon et al. (2023). By default, inputs consist of 4096 tokens and outputs of 1024 tokens. We report the averaged inference throughput (tokens/second) from 5 repeated runs. Unless otherwise specified, all experiments are conducted on NVIDIA Ampere A100 GPUs (40GB).
LLM Evaluation Setup.
Following prior works Biderman et al. (2023); Zhang et al. (2024), we evaluate pretrained models in the zero-shot setting using lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness on nine benchmarks: ARC-Easy Clark et al. (2018), ARC-Challenge Clark et al. (2018), LAMBADA Paperno et al. (2016), HellaSwag Zellers et al. (2019), OpenBookQA Mihaylov et al. (2018), PIQA Bisk et al. (2020), SciQ Welbl et al. (2017), WinoGrande Sakaguchi et al. (2021), and CoQA Reddy et al. (2019).
Fitting Scaling Laws.
Following Gadre et al. (2024); Bian et al. (2025), we use the Levenberg-Marquardt algorithm to fit the conditional scaling laws (Eq. 3). The Levenberg–Marquardt algorithm does least-squares curve fitting by estimating $\hat{\beta}$ as the solution to $\arg\min_{\beta}\sum_{i=1}^{m}\left[y_{i}-f(x_{i},\beta)\right]^{2}$, where $(x_{i},y_{i})$ are the observed data pairs. Note that instead of fitting the Chinchilla scaling law, we empirically searched over architecture variants to find the optimal loss $L_{\text{opt}}(N,D)$ for $N_{\text{non-embed}}<$1B scale.
We scale up the scale law fitting in the following progressive manner:
  •  (Task 1)
    fit on the 8080M results and evaluate on 145145M results;
  •  (Task 2)
    fit on 80,14580,145M results and evaluate on 297297M results;
  •  (Task 3)
    fit on 80,145,29780,145,297M results and evaluate on 11B results;
This ensures a robust and consistent way of scaling up the model sizes and evaluating our conditional scaling law. Following prior work Kumar et al. (2024), we evaluate the fitted scaling law with mean squared error (MSE) metric, defined as $\frac{1}{n}\sum_{i=1}^{n}(l_{i}-\hat{l}_{i})^{2}$ where $l_{i}$ denotes the actual loss and $\hat{l}_{i}$ the predicted loss. We additionally report the Spearman’s rank correlation coefficient Spearman (1961) to compare predicted and actual rankings. Both metrics are calculated on the val data points.
Figure 5: Predictive performances of the fitted conditional scaling law on: (left) Task 1: Fit on 8080M, evaluate on 145145M; (center) Task 2: Fit on 80,14580,145M, evaluate on 297297M; (right) Task 3: Fit on 80,145,29780,145,297M, evaluate on 11B. Orange dots denote fitting data points, and purple crosses indicate the test data points. We compare scaling-law predicted loss with actual pretraining loss of architectures and observed a consistently low MSE and high Spearman correlation across model scales.
5 Experiment Results
We begin by evaluating the predictive performances of the conditional scaling laws with multiplicative calibration. We then conduct ablation studies to assess the impact of data selection and to evaluate the performance of the scaling laws under additive calibration. Finally, we apply the fitted scaling laws to guide the training of large-scale models following the search framework (§5.1).
Predictive Accuracy.
As Task 1-3 described in §4, we fit the conditional scaling laws on 8080M, (8080M, 145145M), and (8080M, 145145M, 297297M) loss-architecture data points, and subsequently evaluate on 145145M, 297297M, and 11B data, respectively. In Figure 5, the low MSE and high Spearman correlation in tasks across different model scales validate the effectiveness and strong predictive performance of the proposed conditional scaling laws.
Ablation of Outliers.
The mlp-to-attention ratio $r_{\text{mlp}/\text{attn}}$ of open-weights models typically fall between $0.5$ and $5$, for example, the mlp-to-attention ratio for LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen3-8B are 4.81, 1.5, and 4.67, respectively. In Figure 5, we fit the conditional scaling law using only model architectures with $r_{\text{mlp/attn}}\in[0.5,5]$. We ablate this choice by training model architectures with outlier $r_{\text{mlp}/\text{attn}}$ below $0.5$ and above $5$ (such as $0.1,12.6$) in Appendix C. In Figure 15 (left) and Figure 15 (center) in Appendix G, we show on Task 3 a comparison of fitting the conditional scaling law without and with these outliers (with a clear Spearman correlation score degradation), which suggests to exclude extreme outliers for better predicted performances.
Ablation of Calibration.
In Figure 15 (right), We ablate an alternative formulation of the scaling laws with additive calibration, as discussed in §3.3. The results on Task 3 show that multiplicative and additive calibrations achieve similar MSE and Spearman correlations, underscoring the robustness of our two-step reference plus calibration framework.
Table 1: Large-Scale Model Results: We evaluate the scaling laws and framework at the 1B and 3B scales by training Panda-1B, Surefire-1B, and Panda-3B, and compare them with LLaMA-3.2-1B and LLaMA-3.2-3B, respectively. The Avg. column reports the mean accuracy across the nine downstream tasks. Panda-1B and Panda-3B are trained using the optimal architectural configurations predicted by our scaling laws, whereas Surefire-1B and Surefire-3B satisfy the loss constraint in Eq. (4) and achieve Pareto optimality.
Models $d_{\text{model}}$ $f_{\text{size}}$ $n_{\text{layers}}$ GQA $d_{\text{model}}/\sqrt{N}$ $r$ Loss ($\downarrow$) Avg. ($\uparrow$)
LLaMA-3.2-1B 2048 8192 16 4 0.066 4.80 2.803 54.9
Panda-1B 2560 4096 16 4 0.082 1.07 2.782 57.0
Surefire-1B 2560 6144 16 9 0.082 3.6 2.804 55.4
LLaMA-3.2-3B 3072 8192 28 3 0.058 4.80 2.625 61.9
Panda-3B 4096 4096 28 3 0.077 1 2.619 62.5
Surefire-3B 4096 4096 28 7 0.077 1 2.620 62.6
Figure 6: Results for 1B and 3B models: (left) Panda-1B closely follows the scaling law predictions for minimizing training loss. (center) Inference throughput comparison between LLaMA-3.2-1B and Surefire-1B, showing that Surefire-1B consistently achieves higher efficiency across batch sizes. (right) Inference throughput comparison between LLaMA-3.2-3B and Surefire-3B, demonstrating that Surefire-3B consistently delivers higher efficiency across all batch sizes.
5.1 Optimal Model Architecture
Validating the conditional scaling law.
We validate the conditional scaling law at the 1B scale by applying multiplicative calibration on Task 3 using data from the (8080M, 145145M, and 297297M) model variants. The learned parameters are
$a_{0}=2.697,a_{1}=0.0974,a_{2}=0.0078,b_{0}=0.3870,b_{1}=0.0063,\text{and}~b_{2}=0.0065.$
From this, we obtain the optimal architectural configuration of $d_{\text{model}}/\sqrt{N}=0.08,r=1.032$ for 1B model by solving $\frac{\partial L}{\partial d_{\text{model}}}=0$ and $\frac{\partial L}{\partial r}=0$. Using this configuration, we train a LLaMA-3.2-style 11B dense model on 100B tokens, denoted as Panda-11B. Panda-11B outperforms the open-weight LLaMA-3.2-11B baseline configs by 2.1% on average across downstream tasks (Table 1). Figure 6 (left) further confirms the effectiveness of the conditional scaling law by showing that Panda-11B achieves the lowest training loss among the exhaustively trained 11B variants under the same setup.
We also scale up our methodology to 3B models. Using the same approach but with data from the 8080M, 145145M, 297297M, and 11B variants, we fit the scaling law and obtain $d_{\text{model}}/\sqrt{N}=0.08$ and $r=1.055$ for the Panda 3B model. Trained on 100B tokens, Panda-3B outperforms the open weight LLaMA-3.2-3B configuration by 0.6% on average across downstream tasks (Table 1).
With all components in place, we apply the search framework for inference-efficient and accurate models (Alg. 1). For the $N_{\text{non-embed}}=1$B and 33B setting trained on 100100B tokens, we set the target loss $L_{t}$ to match the training loss achieved by the LLaMA-3.2-1B and LLaMA-3.2-3B architectures, respectively. Although inference efficiency $I_{N}(P)$ could, in principle, be expressed analytically, it depends heavily on hardware and inference configurations. Therefore, rather than solving for $I_{N}(P)$ directly, we search over feasible configurations $P_{i}$} that satisfy the loss constraint and select Pareto-optimal points, which we denote as Surefire-1B and Surefire-3B. Surefire-1B and Surefire-3B outperform LLaMA-3.2-1B and LLaMA-3.2-3B on downstream tasks (Table 1) and deliver up to 42% higher inference throughput (Figure 6, center and right). Detailed downstream task accuracies are provided in Appendix I.
Table 2: 3B Model Ablation Study: We evaluate the robustness of scaling laws at 3B scale by training Panda-3B and Panda-3B, and compare them with LLaMA-3.2-3B. The Avg. column reports the mean accuracy across the nine downstream tasks. Panda-3B represents the optimal architectural configuration predicted by the conditional scaling laws fitted using the 80M, 145M, and 297M model data, whereas Panda-3B corresponds to the optimal configuration predicted from scaling laws fitted using the 1B model data.
Models $d_{\text{model}}$ $f_{\text{size}}$ $n_{\text{layers}}$ GQA $d_{\text{model}}/\sqrt{N}$ $r$ Loss ($\downarrow$) Avg. ($\uparrow$)
LLaMA-3.2-3B 3072 8192 28 3 0.058 4.80 2.625 61.9
Panda-3B 4096 4096 28 3 0.077 1 2.619 62.5
Panda-3B 4096 4608 28 3 0.076 1.23 2.606 62.5
Figure 7: Effect of the Fitting Dataset on Predictive Performance vs (left) Fit on 80, 145, 297M, 1B, evaluate on 3B; (right) Fit on 1B, evaluate on 3B. Orange dots denote fitting data points, and purple crosses indicate the test data points. We compare scaling-law predicted loss with actual pretraining loss of architectures and we observe that fitting the scaling laws with only 1B model data yields lower MSE and higher Spearman correlation for the 3B model loss prediction.
Ablation of fitting data. While we adopt a progressive strategy for selecting fitting data across tasks (§4), results from small models (e.g., 80M) may not reliably predict behavior at larger scales such as 3B. To examine this, we perform an ablation study by fitting the conditional scaling law for the 3B model using only results from the 1B variant. As shown in Figure 7, fitting with only the 1B data achieves a lower MSE and a higher Spearman correlation when predicting the 3B loss. This indicates that the coefficients of the conditional scaling law may shift as model size increases.
We therefore refit the law with multiplicative calibration using only the 1B variants, yielding the coefficients
$a_{0}=2.319,a_{1}=0.238,a_{2}=0.0176,b_{0}=0.5104,b_{1}=0.0051,\text{and}~b_{2}=0.0062.$
This produces an alternative optimal configuration for the 3B model, with $d_{\text{model}}/\sqrt{N}=0.074$ and $r=1.229$. We train a 3B model (Panda-3B) under this configuration on 100B tokens and compare it with both LLaMA-3.2-3B and Panda-3B (fitted from 80M, 145M, 297M, and 1B data). As shown in Table 2, Panda-3B achieves a lower training loss and comparable downstream accuracy to Panda-3B, with detailed results given in Appendix I. These findings suggest that when scaling up, it is often sufficient, and sometimes preferable, to fit the law using models within a closer size range to the target, such as about one third of its scale.
6 Related Work
Large Language Models.
Transformers Vaswani (2017) have shown strong performance across diverse downstream tasks, such as text classification Wang (2018); Sarlin et al. (2020), mathematical reasoning Cobbe et al. (2021); Hendrycks et al. (2021), and code generation Chen et al. (2021); Austin et al. (2021); Jain et al. (2024). The Transformer architecture serves as the foundation for many leading large language models, including GPT Brown et al. (2020); Achiam et al. (2023), LLaMA Touvron et al. (2023), Gemma Team et al. (2024a), Qwen Yang et al. (2025), Kimi Team et al. (2025), and DeepSeek Liu et al. (2024a); Guo et al. (2025).
Scaling Laws for Language Models.
Scaling laws are powerful tools to predict the performance of large language models. Existing scaling laws Hoffmann et al. (2022); Muennighoff et al. (2023); Sardana et al. (2023); Kumar et al. (2024); Gadre et al. (2024); Ruan et al. (2024) characterize how model performance varies with model size, dataset size, data quality, and compute budget. With the rise of Mixture-of-Experts (MoE) Shazeer et al. (2017); Guo et al. (2025), a powerful architecture for large language models, recent studies Krajewski et al. (2024); Abnar et al. (2025) extend scaling laws to account for the number of experts, expert granularity, active parameters, and sparsity.
Serving Systems.
Due to the increased inference cost, many inference systems have been developed to speed up model serving Yu et al. (2022); Kwon et al. (2023); Zheng et al. (2023); Ye et al. (2025). Specifically, vLLM Kwon et al. (2023) proposes PagedAttention to manage KV cache memory more effectively, thereby improving throughput. Similarly, SGLang Zheng et al. (2023) introduces RadixAttention to achieve higher throughput and lower latency.
Inference-Efficient Model Design.
Efforts to improve the inference efficiency of large language models generally fall into two categories: one line of work investigates the trade-offs across different model configurations Alabdulmohsin et al. (2023); Bian et al. (2025), while the other focuses on designing more efficient model architectures Xiao et al. (2023); Gu & Dao (2023); Gao et al. (2024b); Jiang et al. (2024); Liu et al. (2024b); Dao & Gu (2024); Xiao et al. (2024); Yuan et al. (2025); Chandrasegaran et al. (2025).
7 Limitations and Future Work
While our team has made notable progress, several open challenges remain that offer promising directions for future research. First, due to limitations in resources and time, our evaluation does not extend to 7B models. Second, our analysis is restricted to dense models, and it remains unclear whether the results extend to Mixture of Experts (MoE) architectures Shazeer et al. (2017). While we report inference efficiency measurements for MoE models under varying architectural choices in Appendix J, we have not yet established scaling laws for MoE architectures. Third, we adopt the experimental setup from Chen et al. (2025), and it is uncertain whether different model architectures warrant different hyperparameter configurations. Finally, our analysis is limited to pre-training, and it remains unclear how the results would change under post-training.
8 Conclusion
This work explores the trade-off between model accuracy and inference cost under a fixed training budget. We begin by demonstrating how architectural choices influence both inference throughput and model accuracy. Building on this, we extend Chinchilla scaling laws to incorporate architectural factors and propose a framework for optimal model architecture search. Using the fitted scaling laws and our framework, we trained models up to 3B parameters, achieving up to 42% higher inference throughput and 2.1% accuracy gains across nine downstream tasks.
Reproducibility Statement
All experiments in this work were conducted using publicly available frameworks. Section 4 provides details of our training, inference, and evaluation setups. In particular, we used Megatron-LM (Shoeybi et al., 2019) for model training, vLLM (Kwon et al., 2023) for efficient inference, and lm-eval-harness (Gao et al., 2024a) for standardized evaluations. To facilitate reproducibility, we will release configuration files and scripts.
Appendix A LLM Usage
We used an LLM to improve the writing by correcting grammar in our draft. It was not used to generate research ideas.
Appendix B Open-Weighted Model Architectures
Table 3 presents an overview of the open-weight model architectures utilized in this paper.
Table 3: Open-Weighted Model Architectures: We list the architectural configurations of all models used in this paper. $n_{\text{layers}}$ is the number of layers, $d_{\text{model}}$ is the hidden size, $n_{\text{heads}}$ is the number of attention heads, and $f_{\text{size}}$ is the intermediate size.
Model Name $n_{\text{layers}}$ $d_{\text{model}}$ $n_{\text{heads}}$ $f_{\text{size}}$ GQA
Qwen2.5-1.5B 28 1536 12 8960 6
Qwen3-0.6B 28 1024 16 3072 2
Appendix C Model Architectures
Table LABEL:tab:model_arch provides an overview of the model architectures, all configured with GQA = 4 and employing LLaMA-3.2 as the tokenizer.
Table 4: Model Architectures: We list the architectural configurations of all models trained in this paper. $N_{\text{non-embed}}$ is the total number of non-embedding parameters, $n_{\text{layers}}$ is the number of layers, $d_{\text{model}}$ is the hidden size, $n_{\text{heads}}$ is the number of attention heads, $f_{\text{size}}$ is the intermediate size, and $r_{\text{mlp}/\text{attn}}$ is the MLP-to-attention ratio.
$N_{\text{non-embed}}$ Variant $n_{\text{layers}}$ $d_{\text{model}}$ $n_{\text{heads}}$ $f_{\text{size}}$ $d_{\text{model}}/\sqrt{N}$ $r_{\text{mlp}/\text{attn}}$
80M v1 12 768 16 2048 0.086 2.40
80M v2 12 768 4 2688 0.086 12.6
80M v3 12 768 8 2560 0.085 6.00
80M v4 12 768 24 1536 0.087 1.20
80M v5 12 768 32 1152 0.086 0.68
80M v6 12 768 40 768 0.086 0.36
80M v7 12 768 48 256 0.087 0.10
80M v8 12 384 32 4096 0.043 2.40
80M v9 12 384 8 5376 0.043 12.6
80M v10 12 384 16 5120 0.042 6.00
80M v11 12 384 48 3072 0.044 1.20
80M v12 12 384 64 2304 0.043 0.68
80M v13 12 384 80 1536 0.043 0.36
80M v14 12 384 96 512 0.044 0.10
80M v15 12 1536 8 1024 0.171 2.40
80M v16 12 1536 4 1280 0.169 6.00
80M v17 12 1536 12 768 0.174 1.20
80M v18 12 1536 16 640 0.169 0.75
80M v19 12 1536 20 384 0.171 0.36
80M v20 12 1536 24 128 0.174 0.10
80M v21 12 512 24 3072 0.057 2.40
80M v22 12 512 12 3840 0.056 6.00
80M v23 12 512 16 3584 0.057 4.20
80M v24 12 512 36 2304 0.058 1.20
80M v25 12 512 48 1792 0.057 0.70
80M v26 12 512 60 1152 0.057 0.36
80M v27 12 512 72 384 0.058 0.10
80M v28 12 1024 12 1536 0.114 2.40
80M v29 12 1024 8 1792 0.113 4.20
80M v30 12 1024 16 1280 0.115 1.50
80M v31 12 1024 24 896 0.114 0.70
80M v32 12 1024 36 256 0.114 0.13
80M v33 12 2048 4 896 0.226 4.20
80M v34 12 2048 8 640 0.231 1.50
80M v35 12 2048 16 256 0.226 0.30
80M v48 12 768 20 1792 0.086 1.68
80M v49 12 768 28 1408 0.086 0.94
80M v50 12 384 40 3584 0.043 1.68
80M v51 12 384 52 3072 0.043 1.11
80M v52 12 384 56 2816 0.043 0.94
80M v53 12 384 60 2560 0.043 0.80
80M v54 12 512 32 2560 0.058 1.50
80M v55 12 512 40 2176 0.057 1.02
80M v56 12 512 44 1920 0.058 0.82
80M v57 12 1024 20 1152 0.113 1.08
145M v1 12 1024 16 3072 0.085 3.60
145M v2 12 1024 8 3584 0.084 8.40
145M v3 12 1024 24 2560 0.086 2.00
145M v4 12 1024 32 2304 0.084 1.35
145M v5 12 1024 40 1792 0.085 0.84
145M v6 12 1024 48 1280 0.086 0.50
145M v7 12 1024 64 512 0.085 0.15
145M v8 12 512 32 6144 0.043 3.60
145M v9 12 512 16 7168 0.042 8.40
145M v10 12 512 48 5120 0.043 2.00
145M v11 12 512 64 4608 0.042 1.35
145M v12 12 512 80 3584 0.043 0.84
145M v13 12 512 96 2560 0.043 0.50
145M v14 12 512 128 1024 0.043 0.15
145M v15 12 2048 8 1536 0.170 3.60
145M v16 12 2048 4 1792 0.168 8.40
145M v17 12 2048 12 1280 0.172 2.00
145M v18 12 2048 16 1152 0.168 1.35
145M v19 12 2048 20 896 0.170 0.84
145M v20 12 2048 24 640 0.172 0.50
145M v21 12 2048 32 256 0.170 0.15
145M v22 12 768 24 3840 0.065 3.00
145M v23 12 768 32 3584 0.063 2.10
145M v24 12 768 40 3072 0.064 1.44
145M v25 12 768 48 2560 0.065 1.00
145M v26 12 768 56 2304 0.063 0.77
145M v27 12 768 64 1792 0.064 0.53
145M v28 12 1536 12 1920 0.129 3.00
145M v29 12 1536 16 1792 0.127 2.10
145M v30 12 1536 20 1536 0.128 1.44
145M v31 12 1536 24 1280 0.129 1.00
145M v32 12 1536 28 1152 0.127 0.77
145M v33 12 1536 32 896 0.128 0.53
145M v34 12 4096 4 768 0.340 3.60
145M v35 12 4096 16 128 0.340 0.15
145M v48 12 1024 28 2368 0.086 1.59
145M v49 12 1024 36 2048 0.085 1.07
145M v50 12 512 52 5120 0.042 1.85
145M v51 12 512 60 4800 0.042 1.50
145M v52 12 512 68 4224 0.043 1.16
145M v53 12 512 72 3968 0.043 1.03
145M v54 12 768 44 2944 0.063 1.25
145M v55 12 768 52 2432 0.064 0.88
297M v1 12 1536 24 4096 0.089 3.20
297M v2 12 1536 8 4864 0.090 11.4
297M v3 12 1536 16 4608 0.088 5.40
297M v4 12 1536 32 3584 0.090 2.10
297M v5 12 1536 48 2816 0.089 1.10
297M v6 12 1536 64 2048 0.088 0.60
297M v7 12 1536 80 1024 0.090 0.24
297M v8 12 768 48 8192 0.045 3.20
297M v9 12 768 16 9728 0.045 11.4
297M v10 12 768 32 9216 0.044 5.40
297M v11 12 768 64 7168 0.045 2.10
297M v12 12 768 96 5632 0.045 1.10
297M v13 12 768 128 4096 0.044 0.60
297M v14 12 768 160 2048 0.045 0.24
297M v15 12 3072 12 2048 0.178 3.20
297M v16 12 3072 4 2432 0.180 11.4
297M v17 12 3072 8 2304 0.177 5.40
297M v18 12 3072 16 1792 0.180 2.10
297M v19 12 3072 24 1408 0.178 1.10
297M v20 12 3072 32 1024 0.177 0.60
297M v21 12 3072 40 512 0.180 0.24
297M v22 12 1024 36 6144 0.059 3.20
297M v23 12 1024 12 7296 0.060 11.4
297M v24 12 1024 24 6912 0.059 5.40
297M v25 12 1024 48 5376 0.060 2.10
297M v26 12 1024 72 4224 0.059 1.10
297M v27 12 1024 96 3072 0.059 0.60
297M v28 12 1024 120 1536 0.060 0.24
297M v29 12 2048 12 3456 0.118 5.40
297M v30 12 2048 24 2688 0.120 2.10
297M v31 12 2048 48 1536 0.118 0.60
297M v32 12 2048 60 768 0.120 0.24
297M v45 12 1536 40 3200 0.089 1.50
297M v46 12 1536 44 3072 0.089 1.31
297M v47 12 1536 52 2688 0.088 0.97
297M v48 12 1536 56 2432 0.089 0.81
297M v49 12 768 80 6400 0.045 1.50
297M v50 12 768 88 6016 0.045 1.28
297M v51 12 768 104 5376 0.044 0.97
297M v52 12 768 112 4736 0.045 0.79
297M v53 12 3072 20 1664 0.177 1.56
297M v54 12 3072 28 1152 0.180 0.77
297M v55 12 1024 56 4864 0.060 1.63
297M v56 12 1024 64 4608 0.060 1.35
297M v57 12 1024 80 3840 0.059 0.90
297M v58 12 1024 88 3328 0.060 0.71
297M v59 12 2048 32 2432 0.117 1.43
297M v60 12 2048 36 2048 0.120 1.07
297M v61 12 2048 40 1920 0.118 0.90
297M v62 12 2048 44 1792 0.117 0.76
1B v1 16 2048 32 8192 0.066 4.80
1B v2 16 2048 72 5760 0.067 1.50
1B v3 16 2816 92 2432 0.089 0.50
1B v4 16 2816 76 3072 0.091 0.76
1B v5 16 2816 68 3584 0.090 0.99
1B v6 16 2816 60 4096 0.090 1.28
1B v7 16 2816 56 4480 0.089 1.50
1B v8 16 2816 24 6144 0.089 4.80
1B v9 16 2816 48 4736 0.090 1.85
1B v10 16 2816 40 5120 0.090 2.40
1B v11 16 2816 36 5376 0.090 2.80
1B v12 16 2560 64 4480 0.082 1.31
1B v13 16 2560 72 4096 0.082 1.07
1B v14 16 2560 80 3648 0.082 0.86
1B v15 16 2560 56 4864 0.082 1.63
1B v16 16 2560 88 3200 0.082 0.68
1B v17 16 2560 48 5376 0.082 2.10
Appendix D Hyper-parameters
Table 5 lists the detailed hyper-parameters used for training in this paper.
Table 5: Hyper-parameters: We show the hyper-parameters used for training in this paper.
Model Size 80M 145M 297M 1B 3B
Batch Size 256 256 512 512 512
Max LR 1.5e-3 1.0e-3 8.0e-4 6.0e-4 6.0e-4
Min LR $0.1\times$ Max LR
Optimizer AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.95$)
Weight Decay $0.1$
Clip Grad Norm $1.0$
LR Schedule Cosine
Warmup Steps $500$
Sequence Length $2048$
Appendix E Additional Inference Evaluation Results
In this section, we present additional inference efficiency results on NVIDIA A100 GPUs. Figure 10 presents that, when parameter count, MLP-to-Attention ratio, and hidden size are fixed, increasing GQA leads to higher inference throughput, consistent with the findings of Ainslie et al. (2023). We alter model configurations of LLaMA-3.2-1B, 3B, and LLaMA-3.1-8B in Figure 10.
Figure 8: Hidden size on Inference Throughput: (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. Across varying batch sizes and model scales, larger hidden sizes yield higher inference throughput under a fixed parameter budget. The legend indicates the hidden size of the models, where $d=d_{\text{model}}$.
Figure 9: MLP-to-Attention ratio on Inference Throughput: (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. Across varying batch sizes and model scales, a larger MLP-to-Attention ratio increases inference throughput under a fixed parameter budget. The legend indicates the MLP-to-Attention ratio of the models, where $r=r_{\text{mlp}/\text{attn}}$.
Figure 10: GQA on Inference Throughput: (left) 1B model variants; (center) 3B model variants; (right) 8B model variants. This figure shows the impact of GQA on inference throughput. With the total parameter count fixed, hidden size is set to 2048 (1B), 3072 (3B), and 4096 (8B), and the MLP-to-Attention ratio is 4.0, 2.67, and 4.2, respectively. Across varying batch sizes, models with larger GQA achieve higher throughput. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
Furthermore, we derive architectural variants by altering the configurations of Qwen3-0.6B, 1.7B, and 4B to investigate the impact of model architectural factors on inference efficiency. The results are shown in Figure 11-13.
Figure 11: Hidden size on Inference Throughput (Qwen3): (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, larger hidden sizes yield higher inference throughput under a fixed parameter budget. The legend indicates the hidden size of the models, where $d=d_{\text{model}}$. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
Figure 12: MLP-to-Attention ratio on Inference Throughput (Qwen3): (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. Across varying batch sizes and model scales, a larger MLP-to-Attention ratio increases inference throughput under a fixed parameter budget. The legend indicates the MLP-to-Attention ratio of the models, where $r=r_{\text{mlp}/\text{attn}}$. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
Figure 13: GQA on Inference Throughput (Qwen3): (left) Qwen3-0.6B model variants; (center) Qwen3-1.7B model variants; (right) Qwen3-4B model variants. This figure shows the impact of GQA on inference throughput. With the total parameter count fixed, hidden size is set to 1024 (0.6B), 2048 (1.7B), and 2560 (4B), and the MLP-to-Attention ratio is 1.5, 3.0, and 2.85, respectively. Across varying batch sizes, models with larger GQA achieve higher throughput. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 4096 input and 1024 output tokens.
Appendix F Additional Results: Loss vs. Model Architecture
In this section, we analyze the relationship between training loss and GQA while fixing the number of parameters, hidden size, and MLP-to-Attention ratio. As shown in Figure 14, unlike hidden size and MLP-to-Attention ratio, the relationship between loss and GQA is highly fluctuating.
Figure 14: Loss vs. GQA: (left) 80M model variants; (center) 145M model variants; (right) 297M model variants. Across different model sizes, the relationship between training loss and GQA varies substantially when hidden size and the mlp-to-attention ratio are fixed. The legend denotes the hidden size of each trained model.
Appendix G More Ablation Study
In this section, We first evaluate the impact of outlier data on the fitting of the scaling laws in Figure 15 (left) and Figure 15 (center). Then, we evaluate the fitting performance of multiplicative calibrations and additive calibrations in Figure 15 (left) and Figure 15 (right).
Finally, we evaluate the performance of Joint and non-separable calibrations shown below in Figure 16:
$(a_{0}+a_{1}\log(\frac{dr}{\sqrt{N}})+a_{2}/(\frac{dr}{\sqrt{N}}))\cdot L_{\text{opt}}$
where $d=d_{\text{model}}$, $r=r_{\text{mlp}/\text{attn}}$, and $N=N_{\text{non-embed}}$. In Figure 16, we observe that the performance of joint and non-separable calibrations is significantly worse than that of multiplicative calibration, consistent with our discussion in §3.3.
Figure 15: Ablation Study: (left) use multiplicative calibrations without outliers; (center) use multiplicative calibrations with outliers; (right) use additive calibrations without outliers. The outlier refers to models trained with an mlp-to-attention ratio below 0.5 or above 5. We observe that outlier data points harm the scaling law fit. Moreover, while multiplicative and additive calibrations differ in formulation, their MSE and Spearman values remain nearly identical. Dots denote the data points used for fitting, while crosses indicate the test data points.
Figure 16: Joint and non-separable calibrations: (left) use multiplicative calibrations; (right) use joint and non-separable calibrations. We observe that joint and non-separable calibrations yield higher MSE and lower Spearman scores than multiplicative calibrations, indicating inferior performance. Dots denote the data points used for fitting, while crosses indicate the test data points.
Appendix H Inference FLOPs Analysis
Building on the inference FLOPs analysis from prior work Kaplan et al. (2020), we begin with the following definition:
  • $d_{\text{model}}$: hidden size
  • $f_{\text{size}}$: intermediate (feed-forward) size
  • $n_{\text{layers}}$: number of layers
  • AA: number of query heads
  • $K$: number of key/value heads
  • $d_{h}$: per-head hidden dimension (query and value)
  • $T$: per-head hidden dim the KV length prior to token generation
Based on the above definition, we have $d_{q}=Ad_{h}$ and $d_{kv}=Kd_{h}$. We focus exclusively on non-embedding FLOPs, resulting in:
Attention: QKV and Project
$n_{\text{layers}}(\underbrace{2d_{\text{model}}d_{q}}_{Q}+\underbrace{2d_{\text{model}}d_{kv}}_{K}+\underbrace{2d_{\text{model}}d_{kv}}_{V}+\underbrace{2d_{\text{model}}d_{q}}_{O})$
Attention: Mask
$n_{\text{layers}}(2Td_{q})$
Feedforward:
$n_{\text{layers}}(3\cdot 2d_{\text{model}}f_{\text{size}})$
Total Inference non-embedding FLOPs:
$\text{Total-FLOPs}=n_{\text{layers}}(\underbrace{2d_{\text{model}}d_{q}}_{Q}+\underbrace{2d_{\text{model}}d_{kv}}_{K}+\underbrace{2d_{\text{model}}d_{kv}}_{V}+\underbrace{2d_{\text{model}}d_{q}}_{O}+\underbrace{2Td_{q}}_{qK^{\top}}+\underbrace{3\cdot 2d_{\text{model}}f_{\text{size}}}_{\text{up, gate, down}})$
Since $P_{\text{non-emb}}\approx n_{\text{layers}}(2d_{\text{model}}d_{q}+2d_{\text{model}}d_{kv}+3d_{\text{model}}f_{\text{size}})$. Therefore, $\text{Total-FLOPs}=2P_{\text{non-emb}}+2n_{\text{layers}}Td_{q}$
we adopt the following three approaches to accelerate inference:
  • Increasing the MLP-to-Attention ratio reduces the term $2Td_{q}$, thereby lowering the total FLOPs.
  • Increasing the hidden size reduces the term $2Td_{q}$, thereby lowering the total FLOPs.
Appendix I More Large-scale Training Results
In this section, we first show the detailed result over downstream tasks of large-scale models in Table 6 and Table 7.
Table 6: Detailed Results on Downstream Tasks for 1B Models: In this table, we show detailed results of 1B models over 9 downstream tasks.
Downstream Tasks LLaMA-3.2-1B Panda-1B Surefire-1B
Arc-Easy 58.8 60.9 59.7
Arc-Challenge 29.8 28.9 30.2
LAMBADA 52.8 55.1 52.0
HellaSwag 56.9 58.4 56.6
OpenBookQA 32.0 33.2 32.0
PIQA 73.6 75.2 73.0
SciQ 84.8 87.2 84.9
WinoGrande 57.1 58.6 57.5
COQA 48.7 55.3 52.7
Avg. 54.9 57.0 55.4
Table 7: Detailed Results on Downstream Tasks for 3B Models: In this table, we show detailed results of 3B models over 9 downstream tasks.
Downstream Tasks LLaMA-3.2-3B Panda-3B Surefire-3B Panda-3B
Arc-Easy 66.4 65.5 67.6 66.8
Arc-Challenge 33.3 35.2 33.9 33.3
LAMBADA 60.6 61.8 61.4 61.5
HellaSwag 66.7 66.9 67.0 67.8
OpenBookQA 38.4 38.6 38.6 38.0
PIQA 76.8 76.9 77.4 76.8
SciQ 89.4 91.2 92.1 90.5
WinoGrande 62.5 63.2 60.5 62.7
COQA 63.3 63.4 65.4 64.9
Avg. 61.9 62.5 62.6 62.5
Appendix J MoE Inference
In this section, we examine how the Mixture-of-Experts (MoE) architecture affects inference efficiency. Figure 17 indicates that larger hidden sizes and higher Active-Experts-to-Attention ratios improve the inference throughput of MoE models, consistent with observations in dense models.
Figure 17: Active-Experts-to-Attn on Inference Throughput: (left) 3B-A1.1B model variants; (center) 5.3B-A1.7B model variants; (right) 8.3B-A1.5B model variants. We study the effect of the Active-Experts-to-Attention ratio on inference throughput by fixing the total number of active parameters, setting GQA to 4, and using a batch size of 2048 to reduce MoE inference variance in this figure. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 1024 input and 256 output tokens.